Work out the exterior and interior angles of a regular polygon with 18 sides
Answer by Alan3354(69443)

Work out the exterior and interior angles of a regular polygon with 18 sides
--------
The sum of the Ext angles is 360 degs for all polygons.
360/18 = 20 degs each
------
Int angles = 180 - 20 = 160 degs each.
18*160 = 2880